# Nash CoT: Multi-Path Inference with Preference Equilibrium

### v2

@@ Abstract
Chain of thought (CoT) is a reasoning framework that can enhance the performance of Large Language Models (LLMs) on complex inference tasks. In particular, among various studies related to CoT, multi-path inference stands out as a simple yet effective improvement. However, there is no optimal setting for the number of inference paths. Therefore, we have to increase the number of inference paths to obtain better results, which in turn increases the inference cost. To address this limitation, we can utilize question-related role templates to guide LLMs into relevant roles, thereby increasing the possibility of correct inferences for each path and further reducing dependence on the number of inference paths while improving reasoning accuracy. However, placing LLMs into specific roles may reduce their reasoning diversity and performance on a few tasks where role dependence is low. To alleviate the excessive immersion of the LLM into a specific role, we propose Nash CoT by constructing a competitive system on each path that balances the generation from role-specific LLMs’ and the general LLMs’ generation, thereby ensuring both effective role adoption and diversity in LLM generation further maintaining the performance of multi-path inference while reducing the requirement of the number of inference paths. We evaluate Nash CoT across various inference tasks, including Arabic Reasoning, Commonsense Question Answering, and Symbolic Inference, achieving results that are comparable to or better than those of multi-path CoT with the equal number of inference paths.
@@ 1 Introduction
Figure 1: Demonstrations of Nash Chain-of-Thought (Nash CoT). As shown in this figure, Nash CoT can be divided into three main steps. Step 1 involves bringing the LLM into a template-related role. Step 2 utilizes the role-immersed LLM and LLM under normal conditions to collect model predictions separately. Step 3 filters the responses based on Nash Equilibrium.
@@
However, excessively casting the LLM into a specific role using templates may compromise its robustness (diversity), thereby adversely affecting its performance in solving problems with minimal role dependency. To address this problem, we can introduce a certain degree of randomness into the reasoning process of the LLM. This can mitigate the constraints imposed by excessive role casting on the reasoning performance, ultimately maintaining the role-immersed multi-path CoT’s reasoning performance in these questions. Driven by this motivation, we construct a gaming system on each path, where the LLM in the role-playing context and the LLM in its normal state act as two distinct players. When this gaming system formed by these two players reaches a Nash Equilibrium, the reasoning under that path can balance the preferences of both role-playing with that of the LLM’s normal state. Furthermore, we can choose the response with the most Nash Equilibrium hits through voting to ensure the performance of LLM reasoning.
@@
  • To our knowledge, we are the first to introduce the concept of Nash Equilibrium into multi-path CoT, with the objective of balancing text generation guided by the role-immersed LLM with that produced by the LLM’s default state. This innovation enables us to minimize the number of paths required to achieve good results in multi-path inference, while simultaneously preserving the multi-path CoT’s inherent performance advantages.
@@ 2 Related Work
Training policy with Reinforcement Learning (RL) to reflect preference, termed Reinforcement Learning with Human Feedback (RLFH), is initially introduced by Akrour et al. (2011) and has since undergone consistent improvement and refinement by Cheng et al. (2011); Busa-Fekete et al. (2013); Wilson et al. (2012). This approach has been widely applied to adjust the parameters of LLMs in order to align them with human preferences (Ouyang et al., 2022; Jiang et al., 2023; etc., 2023). Recently, a new approach called Direct Optimizing from Preference (DPO) has been proposed by Rafailov et al. (2023), aiming to directly adjust LLMs to reflect human preferences without requiring a reward model. Additionally, Munos et al. (2023) proposes combining DPO with Nash Equilibrium to ensure the convergence of the last iterated policy. Despite our study also utilizing the concept of Nash Equilibrium in the Preference model, the main difference is that we utilize Nash Equilibrium as a standard to pick up the answer instead of optimizing the LLM’s parameters.
@@ 3 Preference Equilibrium in mini-batch inference
The experimental results of self-consistency indicate that multi-path reasoning can unleash the reasoning capability of LLMs. However, multi-path inference lacks a theoretical foundation for determining the optimal number of inference paths, requiring us to increase the number of paths to ensure performance, and further increase the computational resource consumption. To reduce the number of paths required for multi-path inference, we utilize the concept of Nash Equilibrium to locally construct a binary gaming system in multi-path inference. Specifically, the preference of each valid inference path of the LLM needs to achieve Nash equilibrium with the preferences of the generation guided by the role-immersed LLM. This approach increases the possibility of each path correctly answering the question while maintaining a certain level of robustness (diversity), thereby reducing the number of inference paths required by self-consistency.
@@
Geographer: You are a geographer with a deep understanding of geographical knowledge. Please approach the given problem from the perspective of a geographer.
@@
subsequently, we imply the necessity and existence of Nash Equilibrium in the Preference model (shown in Equation 1).
@@
We first discuss why we should construct a gaming system: To increase the accuracy of each inference path, we can select the most suitable template (example 3) for the given question, immersing the LLM into a template-related role. This approach helps LLMs to solve problems more effectively, thereby reducing the number of inference paths needed.
However, this may lead to some issues: If the LLM is excessively immersed in a specific role, it may reduce the diversity of LLM’s generation, and further reduce LLM’s accuracy in solving tasks that have low role dependence. To address these problems, we locally build a bi-player gaming system that the preference of role-immersed LLM (player 1) over the normal status of LLM (player 2) is the pay-off of role-immersed LLM, vice visa. If player 1 and player 2 reach Nash Equilibrium  [^2] Preference Equilibrium leverages the concepts of Nash Equilibrium. Nash equilibrium is proposed by John Nash. It is a concept solution where, assuming each participant knows the equilibrium strategies of the other participants, no participant can benefit by changing their own strategy., then the generation can balance the preference of both player 1 and player 2.
@@
Given any two policys (players) $\pi_{1}$ and $\pi_{2}$ along with the pay-off of $\pi_{1}$ and $\pi_{2}$ as $\mathcal{P}(\pi_{1}\prec\pi_{2})$ and $\mathcal{P}(\pi_{2}\prec\pi_{1})$, we define the Nash Equilibrium of $\pi_{1}$ and $\pi_{2}$ as Preference Equilibrium.
Subsequently, we define the status that player 1 and player 2 reach Nash Equilibrium as Preference Equilibrium (Definition 3.1). Meanwhile, in Theorem 3.1, we prove the existence of Nash Equilibrium in this system. In particular, the strategy of player 1 equal to player 2 is one solution that this system has reached Preference Equilibrium.
Theorem 3.1 (Existence of Preference Equilibrium).
Given any two policy (player) $\pi_{1}$ and $\pi_{2}$ within the gaming system defined in Definition 1, where $\pi\in\Pi$. $\pi_{1}\equiv\pi_{2}$ is a solution where the gaming system reaches Nash Equilibrium.
@@
Theorem 3.1 proves the existence of a Nash Equilibrium between the role-immersed LLM and the normal status of LLM generation. Therefore, when reaching Preference Equilibrium, the preference of generations made by the role-immersed LLM is aligned with those made by the LLM without templates. Meanwhile, the preference for role-immersed LLM generation is much closer to the selected role-immersed context, while those of the normal status LLM are predominantly based on its parameters which are much more diverse than role-immersed LLM generation. Therefore, Preference Equilibrium can balance the requirement of role-immersed LLM generation and the diversity of the LLM generation in each path. In particular, $\pi_{1}\equiv\pi_{2}$ means their outputs are also likely to be equal. This insight forms a fundamental basis for practically implementing Nash CoT.
@@ 3.1 Mini-batch inference with Preference Equilibrium
Subsequently, based on Preference Equilibrium, we conceptualize a mini-batch inference (shown in Step 2 of Figure 1) as a bi-player gaming system. This approach aims to maintain the performance benefits of multi-path inference while retaining some of the inherent randomness (diversity) of standard inference methods. Before introducing mini-batch inference, we first define $x^{t}$ as the template of zero-shot CoT. Meanwhile, we leverage LLMs to select a player template from the candidate set $\{x^{c}_{0},x^{c}_{1},\cdots,x^{c}_{n}\}$ for role-immersed generation, guided by Equation 1.
@@
Based on the mini-batch inference, we further introduce Nash CoT in the next chapter. (In particular, in the current framework, successful inference cannot always be guaranteed. For instance, $y^{*}$ may not always in $[y_{1},y_{2}]$. We will address this problem in the following chapters.)
@@ 4 Nash Chain of Thought (Nash CoT)
When generating candidate answers, the process predominantly involves two types of loops: mini-batch Loops ($\rm n_{\rm mini}$): As shown in Algorithm 1, this process involves searching for role-immersed generations within two rounds of generation $[y_{1},y_{2}]$. We denote the times of these two predictions as the $\rm n_{\rm mini}$. Moreover, to mitigate the impact brought by low-frequency predictions, we introduce iterating $\rm n_{\rm mini}$ multiple times. This leads us to another type of loop: Outer Loops ($\rm n_{\rm outer}$): This loop resembles the concept of multi-path inference that iterates $\rm n_{\rm mini}$ multiple times. After completing loop $\rm n_{\rm outer}$, we filter the generated answers and retain the answer that reaches Preference Equilibrium most frequently (shown in Algorithm 2), as the predicted answer.
Core LLM Methods SingleEQ AddSub MultiArith GSM8K AQuA SVAMP Avg.
@@
In terms of answer filtering, as shown in Algorithm 2 we first count the most frequent prediction to satisfy Preference equilibrium. Specifically, we count all $y^{*}$ satisfy $y^{*}\in[y_{1},y_{2}]$ and compute their frequency. Subsequently, we return the most frequent answer. Otherwise, if no cases satisfy $y^{*}\in[y_{1},y_{2}]$, we adapt answer filtering by selecting the most frequent prediction among all generated answers.
@@ 5 Experiments
To validate that Nash CoT is a general CoT method, we selected different large models as test models, including Mistral-Instruct (7B) (Jiang et al., 2023), Chat-GLM4 (9B) Zeng et al. (2022); Du et al. (2022). In particular, all of these selected LLMs are turned via RLHF, and the difference between LLM turned with RLHF and the original foundation models have been detailed by Ouyang et al..
@@
Specifically, in Table 12 and 3, we set the total number of reasoning paths for Nash CoT to 10 and for self-consistency to 20. The primary goal is to demonstrate that Nash CoT can achieve the same or even better results with only half the number of reasoning paths compared to self-consistency.
@@ 6 Ablation Study
We further conduct extensive ablations to answer the following questions: 1) Is there a correlation between the performance of Nash CoT and the setting of the number of inference paths and can the performance of Nash CoT be improved to surpass self-consistency by adjusting the number of inference paths? 2) Does the template improve the accuracy of path predictions, and what impact does it have on experimental performance? 3) Is the Nash Equilibrium the reason for the performance improvement observed in Nash CoT?
@@
Effectiveness of Nash Equilibrium.
While Nash CoT has demonstrated effectiveness across various LLMs and benchmarks, however, the extent to which Nash CoT’s performance is driven by Nash Equilibrium remains to be determined. To further investigate, we compare the self-consistency integrated template (denoted as self-con.+temp (10 paths)) with Nash CoT (10 paths) using Mistral-Instruct (7B). As shown in Table 7, Nash CoT consistently outperforms self-con.+temp when applying the same mathematical templates across selected tasks.
@@
Table 7: Ablation of Nash Equilibrium. This experiment is conducted to verify the necessity of Nash Equilibrium in the performance of Nash CoT.
@@ 7 Conclusion
We prove the existence of a Nash equilibrium in preferences between role-immersed and normal-status LLM generations, and develop a gaming system within each path of the multi-path CoT to introduce Nash CoT. Experimental results show that Nash CoT can perform equally or even better than self-consistency while requiring fewer paths.
@@ Acknowledgement
Ethics Claims
Pre-training or fine-tuning an LLM requires much more computing resources. Multi-path CoT is an ideal approach that has been proven to enhance the performance of LLMs’ inference. Meanwhile, Nash CoT effectively reduces the inference paths needed and times of multi-path inference. We believe our approach can further elevate the effectiveness of multi-path inference, thereby further improving the effectiveness of LLM.
@@ Appendix D Computing Resources
Subsequently, we prove the existence of Nash equilibrium in this system. For any two given polices $\pi_{1}\in\Pi$ and $\pi_{2}\in\Pi$ We first define the pay-off of $\pi_{1}$ and $\pi_{2}$ as $R(\pi_{1};\pi_{2})$ and $R(\pi_{2};\pi_{1})$:
$\begin{split}R(\pi_{1};\pi_{2})=\mathcal{P}(\pi_{1}\prec\pi_{2})\\ R(\pi_{2};\pi_{1})=\mathcal{P}(\pi_{1}\succ\pi_{2}),\end{split}$ (3)
we provide the proof of the existence of Nash equilibrium in this system. We define $\bar{\pi}=[\pi_{1},\pi_{2}]$, $v(\bar{\pi})=[R(\pi_{1};\pi_{2}),R(\pi_{1};\pi_{2})]$. According to the Nash equilibrium, it should have to satisfy this relationship:
$\begin{split}v(\bar{\pi}^{*})(\bar{\pi}^{*}-\bar{\pi})\leq 0\end{split}$ (4)
Subsequently, refer to Munos et al., we can learn that if we want Equation 4 holds true, we just have to guarantee Equation 5 holds true.
$\begin{split}(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})\leq 0,\end{split}$ (5)
where $\bar{\pi}$ and $\bar{\pi}^{\prime}$ are any two given policy set. Subsequently, we can further arrive at the following relationships:
$\begin{split}(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})&=\begin{pmatrix}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\\ R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\\ \end{pmatrix}\cdot\begin{pmatrix}\pi_{1}-\pi_{1}^{\prime},\pi_{2}-\pi_{2}^{\prime}\\ \end{pmatrix}\\ &=\bigg{(}R(\pi_{1};\pi_{2})-R(\pi_{1}^{\prime};\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{(}R(\pi_{2};\pi_{1})-R(\pi_{2}^{\prime};\pi_{1}^{\prime})\bigg{)}\\ &\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime})+\bigg{\{}2-\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})+\\ &\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\bigg{\}}\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}-\pi_{2}+\pi_{2}^{\prime})+2\cdot(\pi_{2}-\pi_{2}^{\prime})\\ &=\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})-2\bigg{)}\cdot(\pi_{2}^{\prime}-\pi_{2})+\\ &\bigg{(}\mathcal{P}(\pi_{1}\prec\pi_{2})-\mathcal{P}(\pi_{1}^{\prime}\prec\pi_{2}^{\prime})\bigg{)}\cdot(\pi_{1}-\pi_{1}^{\prime}).\end{split}$ (6)
In particular, we can find that if $\bar{\pi}\equiv\bar{\pi}^{\prime}$ then $(v(\bar{\pi})-v(\bar{\pi}^{\prime}))^{\rm T}(\bar{\pi}-\bar{\pi}^{\prime})\equiv 0$, thus $\bar{\pi}\equiv\bar{\pi}^{\prime}$ is one solution that $\pi_{1}$ and $\pi_{2}$ has reached equilibrium.
Appendix G Further clarify our settings

### v3

@@ Abstract
Chain of thought (CoT) is a reasoning framework that can enhance the performance of Large Language Models (LLMs) on complex inference tasks. In particular, among various studies related to CoT, multi-path inference stands out as a simple yet effective improvement. However, there is no optimal setting for the number of inference paths. Therefore, we have to increase the number of inference paths to obtain better results, which in turn increases the inference cost. To address this limitation, we can utilize question-related role templates to guide LLMs into relevant roles, thereby increasing the possibility of correct inferences for each path and further reducing dependence on the number of inference paths while improving reasoning accuracy. However, placing LLMs into specific roles may reduce their reasoning diversity and performance on a few tasks where role dependence is low. To alleviate the excessive immersion of the LLM into a specific role, we propose Nash CoT by constructing a game system on each path that balances the generation from role-specific LLMs’ and the general LLMs’ generation, thereby ensuring both effective role adoption and diversity in LLM generation further maintaining the performance of multi-path inference while reducing the requirement of the number of inference paths. We evaluate Nash CoT across various inference tasks, including Arabic Reasoning, Commonsense Question Answering, and Symbolic Inference, achieving results that are comparable to or better than those of multi-path CoT with the equal number of inference paths.
@@ 1 Introduction
Figure 1: Demonstrations of Nash Chain-of-Thought (Nash CoT). As shown in this figure, Nash CoT can be divided into three main steps. Step 1 involves bringing the LLM into a template-related role. Step 2 utilizes the role-immersed LLM and LLM under normal conditions to collect model predictions separately. Step 3 filters the responses to ensure the existence of a unique Nash Equilibrium (NE).
@@
However, excessively casting the LLM into a specific role using templates may compromise its robustness (diversity), thereby adversely affecting its performance in solving problems with minimal role dependency. To address this problem, we can introduce a certain degree of randomness into the reasoning process of the LLM. This can mitigate the constraints imposed by excessive role casting on the reasoning performance, ultimately maintaining the role-immersed multi-path CoT’s reasoning performance in these questions. Driven by this motivation, we construct a game system on each path, where the LLM in the role-playing context and the LLM in its normal state act as two distinct players. When this game system formed by these two players reaches a unique Nash Equilibrium (NE), the reasoning under that path can balance the preferences of both role-playing with that of the LLM’s normal state. Furthermore, we can choose the response with the most NE hits through voting to ensure the performance of LLM reasoning.
@@
  • To our knowledge, we are the first to introduce the concept of NE into multi-path CoT, with the objective of balancing text generation guided by the role-immersed LLM with that produced by the LLM’s default state. This innovation enables us to minimize the number of paths required to achieve good results in multi-path inference, while simultaneously preserving the multi-path CoT’s inherent performance advantages.
@@ 2 Related Work
Training policy with Reinforcement Learning (RL) to reflect preference, termed Reinforcement Learning with Human Feedback (RLFH), is initially introduced by Akrour et al. (2011) and has since undergone consistent improvement and refinement by Cheng et al. (2011); Busa-Fekete et al. (2013); Wilson et al. (2012). This approach has been widely applied to adjust the parameters of LLMs in order to align them with human preferences (Ouyang et al., 2022; Jiang et al., 2023; etc., 2023). Recently, a new approach called Direct Optimizing from Preference (DPO) has been proposed by Rafailov et al. (2023), aiming to directly adjust LLMs to reflect human preferences without requiring a reward model. Additionally, Munos et al. (2023) proposes combining DPO with NE to ensure the convergence of the last iterated policy. Despite our study also utilizing the concept of NE in the Preference model, the main difference is that we utilize NE as a standard to pick up the answer instead of optimizing the LLM’s parameters.
@@ 3 Preference Equilibrium in mini-batch inference
The experimental results of self-consistency indicate that multi-path reasoning can unleash the reasoning capability of LLMs. However, multi-path inference lacks a theoretical foundation for determining the optimal number of inference paths, requiring us to increase the number of paths to ensure performance, and further increase the computational resource consumption. To reduce the number of paths required for multi-path inference, we utilize the concept of NE to locally construct a binary game system in multi-path inference. Specifically, the preference of each valid inference path of the LLM needs to achieve NE with the preferences of the generation guided by the role-immersed LLM. This approach increases the possibility of each path correctly answering the question while maintaining a certain level of robustness (diversity), thereby reducing the number of inference paths required by self-consistency.
@@
subsequently, we imply the necessity and existence of a unique NE in the Preference model.
@@
We first discuss why we should construct a game system: To increase the accuracy of each inference path, we can select the most suitable template (example 3) for the given question, immersing the LLM into a template-related role. This approach helps LLMs to solve problems more effectively, thereby reducing the number of inference paths needed.
However, this may lead to some issues: If the LLM is excessively immersed in a specific role, it may reduce the diversity of LLM’s generation, and further reduce LLM’s accuracy in solving tasks that have low role dependence. To address these problems, we locally build a bi-player game system that the preference of role-immersed LLM (player 1) over the normal status of LLM (player 2) is the pay-off of role-immersed LLM, vice visa. If player 1 and player 2 reach NE, then the generation can balance the preference of both player 1 and player 2.
Regularized form of preference model.
[…]
Unlike the trainable RLHF framework proposed by Munos et al., which aims to regularize the training policy to be close to a known reference (safe) policy $\mu$, our inference system is not trainable. Instead, the incorporation of this regularization term serves a different purpose: Facilitate a theoretical analysis that demonstrates that, under the KL constraint, the game system possesses a unique Nash Equilibrium (NE).
@@
Given any two policy (players) $\pi_{1},\pi_{2}$ along with the pay-off of $\pi_{1},\pi_{2}$ as $\mathcal{P}(\pi_{2}\prec\pi_{1})-\tau\cdot{\rm KL}(\pi_{1}||\mu)$ and $\mathcal{P}(\pi_{1}\prec\pi_{2})-\tau\cdot{\rm KL}(\pi_{2}||\mu)$, we define the NE of $\pi_{1}$, $\pi_{2}$ as Preference Equilibrium.
Subsequently, we define the status that player 1 and player 2 reach NE as Preference Equilibrium (Definition 3.1). Meanwhile, in Theorem 3.1, we prove the uniqueness of NE in this system. Specifically, having the strategy of player 1 equal to player 2 is a crucial prerequisite for ensuring that the game system has a unique Preference Equilibrium.
Theorem 3.1 (Uniqueness of Preference Equilibrium).
Given any two policy (player) $\pi_{1}$ and $\pi_{2}$ within the game system defined in Definition 1, where $\pi\in\Pi$. $\pi_{1}=\pi_{2}$ is an essential condition to guarantee this system has a unique NE.
@@
When reaching Preference Equilibrium, the preference of generations made by the role-immersed LLM is aligned with those made by the LLM without templates. Specifically, the preference for role-immersed LLM generation is much closer to the selected role-immersed context, while those of the normal status LLM are predominantly based on its parameters which are much more diverse than role-immersed LLM generation. Therefore, Preference Equilibrium can balance the requirement of role-immersed LLM generation and the diversity of the LLM generation in each path.
Meanings of $\pi_{1}=\pi_{2}$.
Theorem 3.1 proves $\pi_{1}=\pi_{2}$ is an essential condition that guaranteeing this game system has a unique NE. Therefore, we can choose $\pi_{1}=\pi_{2}$ as the indicator to ensure that this game system has a unique NE.
Meanings of KL terms.
If the payoff of $\pi$ lacks a KL term, proving the uniqueness of the Nash Equilibrium (NE) becomes infeasible. Furthermore, the KL term plays a crucial preferential role in penalizing role-immersed generation, particularly when it diverges significantly from normal generation. This penalty helps to strike a balance between role-immersed and normal generations (see the Proof of Theorem 3.1).
@@ 3.1 Mini-batch inference with Preference Equilibrium
Subsequently, based on Preference Equilibrium, we conceptualize a mini-batch inference (shown in Step 2 of Figure 1) as a bi-player game system. This approach aims to maintain the performance benefits of multi-path inference while retaining some of the inherent randomness (diversity) of standard inference methods. Before introducing mini-batch inference, we first define $x^{t}$ as the template of zero-shot CoT. Meanwhile, we leverage LLMs to select a player template from the candidate set $\{x^{c}_{0},x^{c}_{1},\cdots,x^{c}_{n}\}$ for role-immersed generation, guided by Equation 1.
@@
[…]
@@ 4 Nash Chain of Thought (Nash CoT)
When generating candidate answers, the process predominantly involves two types of loops: mini-batch Loops ($\rm n_{\rm mini}$): As shown in Algorithm 1, this process involves searching for role-immersed generations within two rounds of generation $[y_{1},y_{2}]$. We denote the times of these two predictions as the $\rm n_{\rm mini}$. Moreover, to mitigate the impact brought by low-frequency predictions, we introduce iterating $\rm n_{\rm mini}$ multiple times. This leads us to another type of loop: Outer Loops ($\rm n_{\rm outer}$): This loop resembles the concept of multi-path inference that iterates $\rm n_{\rm mini}$ multiple times. After completing loop $\rm n_{\rm outer}$, We filter the generated answers and retain the one that satisfies the uniqueness of the NE most frequently. (shown in Algorithm 2), as the predicted answer.
@@
Core LLM Methods SingleEQ AddSub MultiArith GSM8K AQuA SVAMP Avg.
@@
In terms of answer filtering, as shown in Algorithm 2 we first count the most frequent prediction to satisfy the uniqueness of the NE. Specifically, we count all $y^{*}$ satisfy $y^{*}\in[y_{1},y_{2}]$ and compute their frequency. Subsequently, we return the most frequent answer. Otherwise, if no cases satisfy $y^{*}\in[y_{1},y_{2}]$, we adapt answer filtering by selecting the most frequent prediction among all generated answers.
@@ 5 Experiments
To validate that Nash CoT is a general CoT method, we selected different large models as test models, including Mistral-Instruct (7B) (Jiang et al., 2023), Chat-GLM4 (9B) Zeng et al. (2022); Du et al. (2022). In particular, all of these selected LLMs are turned via RLHF, and the difference between LLM turned with RLHF and the original foundation models have been detailed by Ouyang et al.
@@
In particular, in Table 12 and 3, we set the total number of reasoning paths for Nash CoT to 10 and for self-consistency to 20. The primary goal is to demonstrate that Nash CoT can achieve the same or even better results with only half the number of reasoning paths compared to self-consistency.
@@ 6 Ablation Study
We further conduct extensive ablations to answer the following questions: 1) Is there a correlation between the performance of Nash CoT and the setting of the number of inference paths and can the performance of Nash CoT be improved to surpass self-consistency by adjusting the number of inference paths? 2) Does the template improve the accuracy of path predictions, and what impact does it have on experimental performance? 3) Is the NE the reason for the performance improvement observed in Nash CoT?
@@
Effectiveness of NE.
While Nash CoT has demonstrated effectiveness across various LLMs and benchmarks, however, the extent to which Nash CoT’s performance is driven by NE remains to be determined. To further investigate, we compare the self-consistency integrated template (denoted as self-con.+temp (10 paths)) with Nash CoT (10 paths) using Mistral-Instruct (7B). As shown in Table 7, Nash CoT consistently outperforms self-con.+temp when applying the same mathematical templates across selected tasks.
@@
Table 7: Ablation of NE. This experiment is conducted to verify the necessity of NE in Nash CoT.
@@ 7 Conclusion
We prove the existence of a NE in preferences between role-immersed and normal-status LLM generations, and develop a game system within each path of the multi-path CoT to introduce Nash CoT. Experimental results show that Nash CoT can perform equally or even better than self-consistency while requiring fewer paths.
@@ Appendix D Computing Resources
Our proof is refereed to Munos et al.
Here, we examine a preference model incorporating a Kullback-Leibler (KL) divergence constraint. Unlike the trainable system proposed by Munos et al., which aims to serve as the policy generating the training data for the preference model or address scenarios requiring adherence to a known safe policy, our inference system is non-trainable. Consequently, a static distance is maintained between the players $(\pi_{1},\pi_{2})$ and a static reference policy $\mu$. The introduction of this regularization term is intended to facilitate a theoretical analysis demonstrating that, under the KL constraint, the game system possesses a unique Nash Equilibrium (NE).
[…]
$\begin{split}\mathcal{P}_{\tau}(\pi_{1}\prec\pi_{2})\overset{def}{=}\mathcal{P}(\pi_{1}\prec\pi_{2})-\tau\cdot{\rm KL}(\pi_{2}(\cdot|x)||\mu(\cdot|x))+\tau\cdot{\rm KL}(\pi_{1}(\cdot|x)||\mu(\cdot|x)).\end{split}$ (3)
Then we prove the existence of NE. Specifically, the mappings $\pi_{1}\mapsto\mathcal{P}(\pi_{1}\prec\pi_{2})$ and $\pi_{2}\mapsto\mathcal{P}(\pi_{1}\prec\pi_{2})$ are linear in $\pi_{1}$ (respectively in $\pi_{2}$) thus $\pi_{1}\mapsto\mathcal{P}_{\tau}(\pi_{1}\prec\pi_{2})$ is concave and $\pi_{2}\mapsto\mathcal{P}_{\tau}(\pi_{1}\prec\pi_{2})$ is convex. Existence of a NE is derived from the mini-max theorem for convex-concave functions Sion (1958).
Next, we substitute Equation 3 into our problem setup. In our setup, model training is not involved, so $\mu$ can be regarded as one of the strategies in $\pi_{1}$ and $\pi_{2}$. Furthermore, we assume $\mu$ to be $\pi_{1}$, and $\pi_{2}$ is the role immersed policy, so we can rewrite the Equation 3 as a two-player game system, with the pay-off of $\pi_{1}$ and $\pi_{2}$:
$\begin{split}R(\pi_{1};\pi_{2})&=\mathcal{P}(\pi_{2}\prec\pi_{1})-\tau\cdot{\rm KL}(\pi_{1}||\mu)\\ &=\mathcal{P}(\pi_{2}\prec\pi_{1})\end{split}$ (4)
and
$\begin{split}R(\pi_{2};\pi_{1})&=\mathcal{P}(\pi_{1}\prec\pi_{2})-\tau\cdot{\rm KL}(\pi_{2}||\mu)\\ &\leq\mathcal{P}(\pi_{1}\prec\pi_{2})\end{split}$ (5)
In particular, when the difference between $\pi_{2}$ and $\mu$ is small enough, R($\pi_{2}$; $\pi_{1}$) can be approximately regarded as $\mathcal{P}(\pi_{1}\prec\pi_{2})$. When the difference between $\pi_{2}$ and $\mu$ is large enough, $\mathcal{P}(\pi_{1}\prec\pi_{2})$ serves as an upper bound for $R(\pi_{2};\pi_{1})$. Therefore, the KL term plays a constraining term in the role-immersed generation process.
Subsequently, to show this game has a unique NE, we have to show the variational inequality is strictly monotone (Rosen, Theorem 2). Meanwhile, followed by the proof of Munos et al. We learn that every NE of the game should satisfy the variational inequality for all $\bar{\pi}$:
$\begin{split}v^{T}(\bar{\pi}^{*})(\bar{\pi}^{*}-\bar{\pi})\leq 0,\end{split}$ (6)
Furthermore, the inequality is strictly monotone if and only if for every $\bar{\pi}_{1}$ and $\bar{\pi}_{2}$ we have that:
$\begin{split}(v(\bar{\pi}_{1})-v(\bar{\pi}_{2}))^{T}(\bar{\pi}_{1}-\bar{\pi}_{2})\leq 0,\end{split}$ (7)
And this inequality has been proved by Munos et al., with equality only at $\bar{\pi}_{1}=\bar{\pi}_{2}$.
Therefore, $\pi_{1}=\pi_{2}$ is an essential condition for maintaining strictly monotone. Meanwhile, strictly monotone serves as the guarantee of the uniqueness of the NE. Therefore, we selected $\pi_{1}=\pi_{2}$ as an indicator to ensure the uniqueness of the NE.
Appendix G Extra Explanation